# The Fourier spectral approach to the spatial discretization of quasilinear hyperbolic systems

### v1

@@ 2.1 Discretization with sharp low-pass filters
Let $s>d/2+1$, $M>0$ and suppose that $\bm{U}^{0}\in H^{s}((2\pi\mathbb{T})^{d})^{n}$ with $\big{\lvert}\bm{U}^{0}\big{\rvert}_{H^{s}}\leq M$. Suppose that for all $j\in\{1,\ldots,d\}$, $A_{j}(\bm{U}_{N})$ satisfies the Assumption A.1 and is self-adjoint. Denote $\bm{U}\in{\mathcal{C}}(I;H^{s}((2\pi\mathbb{T})^{d})^{n})$ the maximal-in-time classical solution to (1.1), and $\bm{U}_{N}\in{\mathcal{C}}(I_{N};H^{s}((2\pi\mathbb{T})^{d})^{n})$ the maximal-in-time classical solution to (2.1). Let $T>0$ be the minimum value of Propositions 2.1 and 2.2. For all ${0\leq t\leq T/\big{\lvert}\bm{U}^{0}\big{\rvert}_{H^{s}}}$, there is a $C>0$, depending only on $s$ and $M$ such that for any $0\leq r\leq s$,
@@
We shall assume that $\bm{U}^{0}\in H^{s+1}((2\pi\mathbb{T})^{d})^{n}$, so that $\bm{U}\in{\mathcal{C}}(I;H^{s+1}((2\pi\mathbb{T})^{d})^{n})\cap{\mathcal{C}}^{1}(I;H^{s}((2\pi\mathbb{T})^{d})^{n})$. The general case $\bm{U}^{0}\in H^{s}((2\pi\mathbb{T})^{d})^{n}$ follows by regularization of the initial data, continuity of the flow map and passing to the limit.
We shall first prove the result for $t\in[0,T/\big{\lvert}\bm{U}^{0}\big{\rvert}_{H^{s}}]$. Recall that $T>0$ is the minimum value of Propositions 2.1 and 2.2, so that we have $\left|\bm{U}(t,\cdot)\right|_{H^{s}}\leq 2\left|\bm{U}^{0}\right|_{H^{s}}$ and $\left|\bm{U}_{N}(t,\cdot)\right|_{H^{s}}\leq 2\left|\bm{U}^{0}\right|_{H^{s}}$ for all $N\in\mathbb{N}$. Denote $\bm{D}_{N}\coloneq\bm{U}-\bm{U}_{N}$ and notice
$\partial_{t}\bm{D}_{N}+\sum_{j=1}^{d}A_{j}(\bm{U}_{N})\partial_{x_{j}}\bm{D}_{N}+\sum_{j=1}^{d}(A_{j}(\bm{U})-A_{j}(\bm{U}_{N}))\partial_{x_{j}}\bm{U}=-(\operatorname{Id}-{\sf S}_{N})\left(\sum_{j=1}^{d}A_{j}(\bm{U}_{N})\partial_{x_{j}}\bm{U}_{N}\right).$
@@
Using integration by parts and the properties of $A_{j}(\bm{U}_{N})$ —Assumption A.1 together with the composition estimate of Proposition A.4 and self-adjointness— and the continuous Sobolev embedding $H^{s}((2\pi\mathbb{T})^{d})^{n}\subset W^{1,\infty}((2\pi\mathbb{T})^{d})^{n}$ (Proposition A.1), the first term on the right-hand side is bounded by $C(\left|\bm{U}_{N}\right|_{H^{s}})\left|\bm{U}_{N}\right|_{H^{s}}\left|{\sf S}_{N}^{1/2}\bm{D}_{N}\right|_{L^{2}}^{2}$. Using the commutator estimate of Proposition A.6 and the composition estimate of Proposition A.4, the second term on the right is bounded by $C(\left|\bm{U}_{N}\right|_{H^{s}})\left|\bm{U}_{N}\right|_{H^{s}}\left|\bm{D}_{N}\right|_{L^{2}}\left|{\sf S}_{N}^{1/2}\bm{D}_{N}\right|_{L^{2}}$. By the composition estimate of Proposition A.4, the third term is bounded by $C(\left|\bm{U}\right|_{H^{s}},\left|\bm{U}_{N}\right|_{H^{s}})\left|\bm{U}\right|_{H^{s}}\left|\bm{D}_{N}\right|_{L^{2}}\left|{\sf S}_{N}^{1/2}\bm{D}_{N}\right|_{L^{2}}$, where we again use Assumption A.1 as well as the boundedness of ${\sf S}_{N}^{1/2}$. In all three estimates, we also used the Cauchy-Schwarz inequality. Altogether, we get
@@
Let us now prove the proposition for general $I^{*}$ compact subset of $I$. Without loss of generality, we can assume $0\in I^{*}$ and we will focus on positive times, $t\in I^{*}\cap\mathbb{R}^{+}$, negative times being obtained by time-symmetry. Let $d/2+1<s<s^{\prime}$ and assume $\bm{U}^{0}\in H^{s^{\prime}}$ (this is for notational convenience, as $s^{\prime}$ will eventually play the role of $s$ in the statement of the proposition). Denote $M^{*}=2\sup_{t\in I^{*}}\left|\bm{U}\right|_{H^{s^{\prime}}}$ and $C^{*}$ the constant depending on $s,s^{\prime},M^{*}$ as in Proposition 2.2, and finally $M^{\prime}=2\left|\bm{U}^{0}\right|_{H^{s^{\prime}}}\exp(C^{*}M^{*}T^{*})$, where $T^{*}\coloneq\sup(I^{*})$. We set
$J_{N}=\big{\{}t\in I_{N}\cap\mathbb{R}^{+}\ :\ \sup_{t^{\prime}\in[0,t]}\left|\bm{U}_{N}(t^{\prime},\cdot)\right|_{H^{s}}\leq M^{*}\text{ and }\sup_{t^{\prime}\in[0,t]}\left|\bm{U}_{N}(t^{\prime},\cdot)\right|_{H^{s^{\prime}}}\leq M^{\prime}\big{\}},$
@@
and our aim is to prove that there exists $N_{0}\in\mathbb{N}$ such that for any $N\geq N_{0}$, $J_{N}\supset I^{*}\cap\mathbb{R}^{+}$. We use the continuity argument. By continuity of $\bm{U}_{N}\in{\mathcal{C}}(I_{N};H^{s^{\prime}}((2\pi\mathbb{T})^{d})^{n})$, we have that $J_{N}$ is a non-empty closed interval. Let us now prove that $J_{N}\cap I^{*}\cap\mathbb{R}^{+}$ is an open subset of $I^{*}\cap\mathbb{R}^{+}$. Let $t\in J_{N}\cap I^{*}\cap\mathbb{R}^{+}$. We can follow the proof of the first part of the proposition replacing $s$ with $s^{\prime}$ and the bound $\big{\lvert}\bm{U}\big{\rvert}_{H^{s}}+\big{\lvert}\bm{U}_{N}\big{\rvert}_{H^{s}}\leq 4\big{\lvert}\bm{U}^{0}\big{\rvert}_{H^{s}}$ with $\big{\lvert}\bm{U}\big{\rvert}_{H^{s^{\prime}}}+\big{\lvert}\bm{U}_{N}\big{\rvert}_{H^{s^{\prime}}}\leq M^{*}+M^{\prime}$ to infer that there exists $C$, depending only on $s^{\prime},M^{*}$ and $T^{*}$ such that for all $0\leq r\leq s^{\prime}$,
$\sup_{t^{\prime}\in[0,t]}\big{\lvert}\bm{U}(t^{\prime},\cdot)-\bm{U}_{N}(t^{\prime},\cdot)\big{\rvert}_{H^{r}}\leq C(M^{*}+M^{\prime})N^{r-s^{\prime}}.$
@@
Applying this estimate with $r=s$, it follows that there exists $N_{0}\in\mathbb{N}$ depending only on $s,s^{\prime},M^{*}$ and $T^{*}$ such that for any $N\geq N_{0}$,
$\sup_{t^{\prime}\in[0,t]}\big{\lvert}\bm{U}_{N}(t^{\prime},\cdot)\big{\rvert}_{H^{s}}\leq\sup_{t^{\prime}\in[0,t]}\big{\lvert}\bm{U}(t^{\prime},\cdot)\big{\rvert}_{H^{s}}+\sup_{t^{\prime}\in[0,t]}\big{\lvert}\bm{U}(t^{\prime},\cdot)-\bm{U}_{N}(t,\cdot)\big{\rvert}_{H^{s}}\leq\frac{2}{3}M^{*}.$
@@
Moreover, using the second part of Proposition 2.2, we have immediately
$\sup_{t^{\prime}\in[0,t]}\big{\lvert}\bm{U}_{N}(t^{\prime},\cdot)\big{\rvert}_{H^{s^{\prime}}}\leq\frac{1}{2}M^{\prime}.$
@@
This shows, using again the continuity of $\bm{U}_{N}\in{\mathcal{C}}(I_{N};H^{s^{\prime}}((2\pi\mathbb{T})^{d})^{n})$, that there exists $\delta>0$ such that ${[t-\delta,t+\delta]\subset J_{N}}$, and hence that $J_{N}\cap I^{*}\cap\mathbb{R}^{+}$ is a non-empty, connected, closed and open subset of $I^{*}\cap\mathbb{R}^{+}$. Hence $J_{N}\supset I^{*}\cap\mathbb{R}^{+}$. Moreover, the desired estimate has been proven along the argument. This concludes the proof. ∎
@@ 2.2 Discretization with smooth low-pass filters
where we used the continuous Sobolev embedding $H^{s}((2\pi\mathbb{T})^{d})^{n}\subset W^{1,\infty}((2\pi\mathbb{T})^{d})^{n}$ (Proposition A.1) and that ${\sf S}_{N}^{1/2}:L^{2}\to L^{2}$ is bounded. Testing the identity (2.7) against $\dot{\bm{U}}_{N}$, using the Cauchy-Schwarz inequality on the right-hand side and inserting the two estimates above yields the desired inequality (2.6). With this estimate in hand, the rest of the proof follows exactly the proof of Proposition 2.2.
@@
The proof follows the proof of Proposition 2.4, with one modification. Consider $\bm{U}_{N}\in{\mathcal{C}}(I_{N};H^{s}((2\pi\mathbb{T})^{d})^{n})$ solution to (2.5) (considering instead $\bm{U}_{N}\in{\mathcal{C}}(I_{N};H^{s}((2\pi\mathbb{T})^{d})^{n})$ solution to (2.4) amounts to replacing $A_{j}^{1}(\bm{U}_{N})$ with $A_{j}(\bm{U}_{N})$ in the right-hand side of the following identity, with no consequence). Then the difference $\bm{D}_{N}\coloneq\bm{U}-\bm{U}_{N}$ satisfies
$\partial_{t}\bm{D}_{N}+\sum_{j=1}^{d}A_{j}(\bm{U}_{N})\partial_{x_{j}}\bm{D}_{N}+\sum_{j=1}^{d}(A_{j}(\bm{U})-A_{j}(\bm{U}_{N}))\partial_{x_{j}}\bm{U}=(\operatorname{Id}-{\sf P}_{N})\left(\sum_{j=1}^{d}A_{j}^{1}(\bm{U}_{N})\partial_{x_{j}}\bm{U}_{N}\right).$
@@ 3 Symmetrizable quasilinear systems
There exists $S(\bm{U})$ a Friedrichs-symmetrizer for the system (1.1), that is, there exists an open set ${\mathcal{U}}\subset\mathbb{R}^{n}$ with $\bm{0}\in{\mathcal{U}}$ such that for all $\bm{U}\in{\mathcal{U}}$, $S(\bm{U})$ is real-valued, symmetric positive definite, and for all $j\in\{1,\ldots,d\}$, $S(\bm{U})A_{j}(\bm{U})$ is symmetric. We assume that all entries of $S(\cdot)$ are polynomial.

### v2

@@ 2.1 Discretization with sharp low-pass filters
Let $s>d/2+1$, $M>0$. Suppose that $\bm{U}^{0}\in H^{s}((2\pi\mathbb{T})^{d})^{n}$ with $\big\lvert\bm{U}^{0}\big\rvert_{H^{s}}\leq M$}. Suppose that for all $j\in\{1,\ldots,d\}$, $A_{j}(\cdot)$ satisfies the Assumption A.1 and is self-adjoint. Denote $\bm{U}\in{\mathcal{C}}(I;H^{s}((2\pi\mathbb{T})^{d})^{n})$ the maximal-in-time classical solution to (1.1), and $\bm{U}_{N}\in{\mathcal{C}}(I_{N};H^{s}((2\pi\mathbb{T})^{d})^{n})$ the maximal-in-time classical solution to (2.1). Let $T>0$ be the minimum value of Propositions 2.1 and 2.2. For all $0\leq t\leq T/\big\lvert\bm{U}^{0}\big\rvert_{H^{s}}$}, there is a $C>0$, depending only on $s$ and $M$ such that for any $0\leq r\leq s$,
@@
We shall first prove the result for $t\in[0,T/\big\lvert\bm{U}^{0}\big\rvert_{H^{s}}]$. Let us assume first that $\bm{U}^{0}\in H^{s+1}((2\pi\mathbb{T})^{d})^{n}$ so that $\bm{U}\in{\mathcal{C}}(I;H^{s+1}((2\pi\mathbb{T})^{d})^{n})\cap{\mathcal{C}}^{1}(I;H^{s}((2\pi\mathbb{T})^{d})^{n})$, the general case being deduced afterwards. Recall that $T>0$ is the minimum value of Propositions 2.1 and 2.2, so that we have $\left|\bm{U}(t,\cdot)\right|_{H^{s}}\leq 2\left|\bm{U}^{0}\right|_{H^{s}}$ and $\left|\bm{U}_{N}(t,\cdot)\right|_{H^{s}}\leq 2\left|\bm{U}^{0}\right|_{H^{s}}$ for all $N\in\mathbb{N}$. Denote $\bm{D}_{N}\coloneq\bm{U}-\bm{U}_{N}$ and notice
$\partial_{t}\bm{D}_{N}+\sum_{j=1}^{d}A_{j}(\bm{U}_{N})\partial_{x_{j}}\bm{D}_{N}+\sum_{j=1}^{d}(A_{j}(\bm{U})-A_{j}(\bm{U}_{N}))\partial_{x_{j}}\bm{U}=-(\operatorname{Id}-{\sf P}_{N})\left(\sum_{j=1}^{d}A_{j}(\bm{U}_{N})\partial_{x_{j}}\bm{U}_{N}\right).$
@@
Using integration by parts and the properties of $A_{j}(\cdot)$ —Assumption A.1 together with the composition estimate of Proposition A.4 and self-adjointness— and the continuous Sobolev embedding $H^{s}((2\pi\mathbb{T})^{d})^{n}\subset W^{1,\infty}((2\pi\mathbb{T})^{d})^{n}$ (Proposition A.1), the first term on the right-hand side is bounded by $C(\left|\bm{U}_{N}\right|_{H^{s}})\left|\bm{U}_{N}\right|_{H^{s}}\left|{\sf S}_{N}^{1/2}\bm{D}_{N}\right|_{L^{2}}^{2}$. Using the commutator estimate of Proposition A.6 and the composition estimate of Proposition A.4, the second term on the right is bounded by $C(\left|\bm{U}_{N}\right|_{H^{s}})\left|\bm{U}_{N}\right|_{H^{s}}\left|\bm{D}_{N}\right|_{L^{2}}\left|{\sf S}_{N}^{1/2}\bm{D}_{N}\right|_{L^{2}}$. By the composition estimate of Proposition A.4, the third term is bounded by $C(\left|\bm{U}\right|_{H^{s}},\left|\bm{U}_{N}\right|_{H^{s}})\left|\bm{U}\right|_{H^{s}}\left|\bm{D}_{N}\right|_{L^{2}}\left|{\sf S}_{N}^{1/2}\bm{D}_{N}\right|_{L^{2}}$, where we again use Assumption A.1 as well as the boundedness of ${\sf S}_{N}^{1/2}$. In all three estimates, we also used the Cauchy–Schwarz inequality. Altogether, we get
@@
Let us now explain why the same result holds in the general case $\bm{U}^{0}\in H^{s}((2\pi\mathbb{T})^{d})^{n}$. Consider $(\bm{U}^{0}_{k})_{k\in\mathbb{N}}$ a sequence (constructed by Fourier truncation) such that for all $k\in\mathbb{N}$, $\bm{U}^{0}_{k}\in H^{s+1}((2\pi\mathbb{T})^{d})^{n}$ and $\bm{U}^{0}_{k}\to\bm{U}^{0}$ in $H^{s}((2\pi\mathbb{T})^{d})^{n}$ as $k\to\infty$, and $\big\lvert\bm{U}^{0}_{k}\big\rvert_{H^{s}}\leq\big\lvert\bm{U}^{0}\big\rvert_{H^{s}}$. Then we can apply the above for each $k\in\mathbb{N}$ and infer that $\bm{U}_{k}$ (respectively $\bm{U}_{k,N}$) the solutions to (1.1) (respectively (2.1)) emerging from the initial data $\bm{U}^{0}_{k}$ satisfy for all $0\leq t\leq T/\big\lvert\bm{U}^{0}\big\rvert_{H^{s}}$} and for any $0\leq r\leq s$,
$\big\lvert(\bm{U}_{k}-\bm{U}_{k,N})(t,\cdot)\big\rvert_{H^{r}}\leq C\,\big\lvert\bm{U}^{0}\big\rvert_{H^{s}}\,N^{r-s},$
where $C>0$ depends only on $s$ and $M$, and in particular is uniform with respect to $k$. We now pass to the limit as $k\to\infty$. By standard estimates on the linearized systems from (1.1) and (2.1) (see e.g. [19, Proposition 7.1.8]), we have $\bm{U}_{k}\to\bm{U}$ and $\bm{U}_{k,N}\to\bm{U}_{N}$ in ${\mathcal{C}}([0,T/\big\lvert\bm{U}^{0}\big\rvert_{H^{s}}];L^{2}((2\pi\mathbb{T})^{d})^{n})$ as $k\to\infty$, where we denote $\bm{U}\in{\mathcal{C}}([0,T/\big\lvert\bm{U}^{0}\big\rvert_{H^{s}}];H^{s}((2\pi\mathbb{T})^{d})^{n})$ (respectively $\bm{U}_{N}\in{\mathcal{C}}([0,T/\big\lvert\bm{U}^{0}\big\rvert_{H^{s}}];H^{s}((2\pi\mathbb{T})^{d})^{n})$) the solutions to (1.1) (respectively (2.1)) emerging from the initial data $\bm{U}^{0}$, as in the statement of the Proposition. Because the above estimate is uniform with respect to $k$, we infer as desired that the limits satisfy for all $0\leq t\leq T/\big\lvert\bm{U}^{0}\big\rvert_{H^{s}}$} and for any $0\leq r\leq s$,
@@
Let us now prove the proposition for general $I^{*}$ compact subset of $I$. Without loss of generality, we can assume $0\in I^{*}$ and we will focus on positive times, $t\in I^{*}\cap\mathbb{R}^{+}$, negative times being obtained by time-symmetry. Let $d/2+1<s_{0}<s$. Denote $M^{*}=2\sup_{t\in I^{*}}\left|\bm{U}\right|_{H^{s_{0}}}$ and $C^{*}$ the constant depending on $s_{0},s,M^{*}$ as in Proposition 2.2 with $s_{0}$ playing the role of $s$ and $s$ playing the role of $s^{\prime}$, and $M=2\left|\bm{U}^{0}\right|_{H^{s}}\exp(C^{*}M^{*}T^{*})$, where $T^{*}\coloneq\sup(I^{*})$. We set
$J_{N}=\big\{t\in I_{N}\cap\mathbb{R}^{+}\ :\ \sup_{t^{\prime}\in[0,t]}\left|\bm{U}_{N}(t^{\prime},\cdot)\right|_{H^{s_{0}}}\leq M^{*}\text{ and }\sup_{t^{\prime}\in[0,t]}\left|\bm{U}_{N}(t^{\prime},\cdot)\right|_{H^{s}}\leq M\big\},$
@@
and our aim is to prove that there exists $N_{0}\in\mathbb{N}$ such that for any $N\geq N_{0}$, $J_{N}\supset I^{*}\cap\mathbb{R}^{+}$. We use the continuity argument. By continuity of $\bm{U}_{N}\in{\mathcal{C}}(I_{N};H^{s}((2\pi\mathbb{T})^{d})^{n})$, we have that $J_{N}$ is a non-empty closed interval. Let us now prove that $J_{N}\cap I^{*}\cap\mathbb{R}^{+}$ is an open subset of $I^{*}\cap\mathbb{R}^{+}$. Let $t\in J_{N}\cap I^{*}\cap\mathbb{R}^{+}$. We can follow the proof of the first part of the proposition replacing the bound $\big\lvert\bm{U}\big\rvert_{H^{s}}+\big\lvert\bm{U}_{N}\big\rvert_{H^{s}}\leq 4\big\lvert\bm{U}^{0}\big\rvert_{H^{s}}$ with $\big\lvert\bm{U}\big\rvert_{H^{s}}+\big\lvert\bm{U}_{N}\big\rvert_{H^{s}}\leq M^{*}+M$ to infer that there exists $C$, depending only on $s,M^{*}$ and $T^{*}$ such that for all $0\leq r\leq s$,
$\sup_{t^{\prime}\in[0,t]}\big\lvert\bm{U}(t^{\prime},\cdot)-\bm{U}_{N}(t^{\prime},\cdot)\big\rvert_{H^{r}}\leq C(M^{*}+M)N^{r-s}.$
@@
Applying this estimate with $r=s_{0}$, it follows that there exists $N_{0}\in\mathbb{N}$ depending only on $s_{0},s,M^{*}$ and $T^{*}$ such that for any $N\geq N_{0}$,
$\sup_{t^{\prime}\in[0,t]}\big\lvert\bm{U}_{N}(t^{\prime},\cdot)\big\rvert_{H^{s_{0}}}\leq\sup_{t^{\prime}\in[0,t]}\big\lvert\bm{U}(t^{\prime},\cdot)\big\rvert_{H^{s_{0}}}+\sup_{t^{\prime}\in[0,t]}\big\lvert\bm{U}(t^{\prime},\cdot)-\bm{U}_{N}(t^{\prime},\cdot)\big\rvert_{H^{s_{0}}}\leq\frac{2}{3}M^{*}.$
@@
Moreover, using the second part of Proposition 2.2 (recall $s_{0}$ playing the role of $s$ and $s$ playing the role of $s^{\prime}$), we have
$\sup_{t^{\prime}\in[0,t]}\big\lvert\bm{U}_{N}(t^{\prime},\cdot)\big\rvert_{H^{s}}\leq\left|\bm{U}^{0}\right|_{H^{s}}\exp(C^{*}M^{*}t)\leq\frac{1}{2}M.$
@@
This shows, using again the continuity of $\bm{U}_{N}\in{\mathcal{C}}(I_{N};H^{s}((2\pi\mathbb{T})^{d})^{n})$, that there exists $\delta>0$ such that $0,t+\delta]\subset J_{N}$}, and hence that $J_{N}\cap I^{*}\cap\mathbb{R}^{+}$ is a non-empty, connected, closed and open subset of $I^{*}\cap\mathbb{R}^{+}$. Hence $J_{N}\supset I^{*}\cap\mathbb{R}^{+}$. Moreover, the desired estimate has been proven along the argument. This concludes the proof. ∎
@@ 2.2 Discretization with smooth low-pass filters
where we used the continuous Sobolev embedding $H^{s}((2\pi\mathbb{T})^{d})^{n}\subset W^{1,\infty}((2\pi\mathbb{T})^{d})^{n}$ (Proposition A.1) and that ${\sf S}_{N}^{1/2}:L^{2}\to L^{2}$ is bounded. Testing the identity (2.7) against $\dot{\bm{U}}_{N}$, using the Cauchy–Schwarz inequality on the right-hand side and inserting the two estimates above yields the desired inequality (2.6). With this estimate in hand, the rest of the proof follows exactly the proof of Proposition 2.2.
@@
The proof follows the proof of Proposition 2.4, with one modification. Consider $\bm{U}_{N}\in{\mathcal{C}}(I_{N};H^{s}((2\pi\mathbb{T})^{d})^{n})$ solution to (2.5) (considering instead the solution to (2.4) amounts to replacing $A_{j}^{1}(\bm{U}_{N})$ with $A_{j}(\bm{U}_{N})$ in the right-hand side of the following identity, with no consequence). Then the difference $\bm{D}_{N}\coloneq\bm{U}-\bm{U}_{N}$ satisfies
$\partial_{t}\bm{D}_{N}+\sum_{j=1}^{d}A_{j}(\bm{U}_{N})\partial_{x_{j}}\bm{D}_{N}+\sum_{j=1}^{d}(A_{j}(\bm{U})-A_{j}(\bm{U}_{N}))\partial_{x_{j}}\bm{U}=(\operatorname{Id}-{\sf S}_{N})\left(\sum_{j=1}^{d}A_{j}^{1}(\bm{U}_{N})\partial_{x_{j}}\bm{U}_{N}\right).$
@@ 3 Symmetrizable quasilinear systems
There exists $S(\cdot)$ a Friedrichs-symmetrizer for the system (1.1), that is, there exists an open set ${\mathcal{U}}\subset\mathbb{R}^{n}$} with $\bm{0}\in{\mathcal{U}}$ such that for all $\bm{U}\in{\mathcal{U}}$, $S(\bm{U})$ is real-valued, symmetric positive definite, and for all $j\in\{1,\ldots,d\}$, $S(\bm{U})A_{j}(\bm{U})$ is symmetric. We assume that all entries of $S(\cdot)$ are polynomial.
Remark 3.1.
As we assumed in Assumption A.1 that entries of $A_{j}(\bm{U})$ are polynomial for all $j\in\{1,\ldots,d\}$, then if Friedrichs-symmetrizers exist it is always possible to select one whose entries are polynomial. Indeed, the assumptions that $S(\cdot)$ is real-valued and symmetric and that for all $j\in\{1,\ldots,d\}$, $S(\cdot)A_{j}(\cdot)$ is symmetric constitutes a system of linear equations for entries of $S$. By considering the corresponding matrix in the field of real rational fractions (since entries of $A_{j}(\bm{U})$ are real polynomials) and performing Gaussian elimination, we see that the system of linear equations can be solved for (non-identically zero) smooth functions if and only if it can be solved for rational fractions, which then can be chosen polynomials after multiplication of all entries by a common multiple of all denominators. Notice however that the domain of hyperbolicity defined as the open set on which $S(\cdot)$ is positive definite depends on the choice of the symmetrizer.